Use the following information for Pleasant View Lodging Company to answer the questions about ratios below Income Statement Sales 650,000 Cost of Goods Sold (383,500) Depreciation (58,500) Earnings before interest and taxes 208,000 Interest expense (65,000) Earnings before taxes 143,000 Taxes (35,750) Net income 107,250 Dividends 21,450 Additions to retained earnings 85,800 Balance Sheet Cash 20,150 Accounts payabl 32,280 Accounts receivable 64,480 Notes Payable 17,934 Inventory 108,810 Total Current Liabilities 50,214 Total Current Assets 193,440 Long-term debt 78,907 Property, Plant, and Equipment 165,230 Total Liabilities 129,121Common stock 32,280 Retained earnings 197.269 Total Owners Equity 229,549 Total Assets 358,670 Total Liabilities and Owners Equity 358,670 Give answers to two decimals What is the current ratio? What is the quick ratio? What is the total asset turnover? What is the inventory turnover? What is the days sales in inventory? What is the receivables turnover? What is the days sales in receivables? What is the total debt ratio? What is the debt-equity ratio?What is the times interest earned ratio? What is the profit margin? What is the return on assets? What is the return on equity?
Answer #1

Current ratio = Current assets / Current liabilities = $193,440 / $50,214 = 3.85

Quick ratio = ( Current assets - Inventory ) / Current liabilities = ( $193,440 - $108,810 ) / $50,214 = 1.69

Total asset turnover = Sales / Total assets = $650,000 / $358,670 = 1.81

Inventory turnover = Cost of goods sold / Inventory = $383,500 / $108,810 = 3.52

Days sales in inventory = ( Inventory / Cost of goods sold ) * Days in year = ( $108,810 / $383,500 ) * 365 days = 103.56 days

Receivables turnover = Sales / Accounts receivables = $650,000 / $64,480 = 10.08

Days sales in receivables = Days in year / Receivables turnover ratio = 365 days / 10.08 = 36.21 days

Total debt ratio = Total liabilities / Total assets = $129,121 / $358,670 = 0.36

Debt-equity ratio = Total liabilities / Total owner's equity = $129,121 / $229,549 = 0.56

Times interest earned ratio = Earnings before interest and taxes / Interest expense = $208,000 / $65,000 = 3.2 times

Profit margin = ( Net income / Sales ) * 100 = ( $107,250 / $650,000 ) * 100 = 16.5%

Return on asset = ( Net income / Total assets) * 100 = ( $107,250 / $358,670 ) * 100 = 29.90%

Return on equity = ( Net income / Total owner's equity ) * 100 = ( $107,250 / $229,549 ) * 100 = 46.72%

Note : In the absence of any information regarding rounding off, all the answers are rounded off to 2 decimal places